OpenStudy (anonymous):

find the vertex line of symmetry and maximum or minimum of f(x). Graph the function f(x)=-5(x+3)^2+3

OpenStudy (zzr0ck3r):

calc? or no clac?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

y` = -10(x+3) y`=0=-10x-30 -10x=30 x = -3 is line of symmetry f(-3) = 3 is critical point to see if its a min or max take second derv -10 thus a max

OpenStudy (zzr0ck3r):

oh

OpenStudy (zzr0ck3r):

do you know how to find vertex?

OpenStudy (anonymous):

NO

OpenStudy (zzr0ck3r):

-5(x+3)^2+3 -5(x^2+6x+9)+3 -5x^2-30x+12 vertex is -b/2a so 30/-10 = -3 f(-3) = 3 this isthe x cordinate of the peak or valley of the paraballa. The fact that the 1st term is negative means it opens down so we know.... x= -3 is the line of symetry and f(-3)=3 or (-3,3) is a maximum

OpenStudy (zzr0ck3r):

calculus was one line faster:)

OpenStudy (anonymous):

Thank you so much for explaining, I am struggling

OpenStudy (zzr0ck3r):

-5(x+3)^2+3 -5(x^2+6x+9)+3 -5x^2-30x-42 vertex is -b/2a so 30/-10 = -3 f(-3) = 3 this isthe x cordinate of the peak or valley of the paraballa. The fact that the 1st term is negative means it opens down so we know.... x= -3 is the line of symetry and f(-3)=3 or (-3,3) is a maximum

OpenStudy (zzr0ck3r):

small typo sorry does not change answer.